3.088 x 10^-19g Ag= ? atoms of Ag
1
Expert's answer
2020-12-06T10:33:30-0500

number of atoms is n*Na = m(Ag)/Ar(Ag)*Na = (3.088*10^-19 g)/(107.8682 g/mol)*(6.02*10^23 1/mol) = 1723 atoms (Na is the Avogadro constant)